# Parameter estimation in fluid flow models from undersampled frequency space data

### v1

@@ 6 Conclusion
We have proposed a new formulation for the inverse problem for parameter estimation in fluid flow problems using undersampled frequency-space MRI data and demonstrated it using a Reduced-Order Unscented Kalman Filter. This method outperforms the results of an inverse problem using velocity data reconstructed with Compressed Sensing, especially for high undersampling rates and different $venc$ values.
@@ Acknowledgments
We acknowledge the funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No 852544 - CardioZoom).
@@ Appendix A Numerical solution method of the forward problem
Given the initial conditions $\bm{u}^{0}=\bm{u}(0)\in V_{\Gamma_{w},h}$ and $\pi^{0}_{1},\dots,\pi_{N}^{0}\in\mathbb{R}$, perform for $j>0$, with $t^{j}=j\tau$:
1. Viscous Step: Find the tentative velocity $\tilde{\bm{u}}^{n}\in V_{\Gamma_{w},h}$ such that:
$\begin{cases}\displaystyle\tilde{\bm{u}}^{j}|_{\Gamma_{in}}=\bm{u}_{inlet}(t^{j})\\ \displaystyle\frac{\tau}{\rho}(\tilde{\bm{u}}^{j},\bm{v})_{\Omega_{h}}+\rho(\bm{u}^{j-1}\cdot\nabla\tilde{\bm{u}}^{j},\bm{v})_{\Omega_{h}}+\frac{\rho}{2}((\nabla\cdot\bm{u}^{j-1})\tilde{\bm{u}}^{j},\bm{v})_{\Omega_{h}}+(\delta\bm{u}^{j-1}\cdot\nabla\tilde{\bm{u}}^{j},\bm{u}^{j-1}\cdot\nabla\bm{v})_{\Omega_{h}}\\ \displaystyle+2\mu(\epsilon(\tilde{\bm{u}}^{j}),\epsilon(\bm{v}))_{\Omega_{h}}+\sum_{\ell=1}^{K}\frac{\rho}{2}|\bm{u}^{j-1}\cdot\mathbf{n}|_{-}(\tilde{\bm{u}}^{j},\bm{v})_{\Gamma_{\ell}}=\frac{\tau}{\rho}(\bm{u}^{j-1},\bm{v})_{\Omega_{h}}\end{cases}$ (20)
@@ 1. Viscous Step: Find the tentative velocity $\tilde{\bm{u}}^{n}\in V_{\Gamma_{w},h}$ such that:
2. Projection-Windkessel Step: Compute $\tilde{Q}^{j}=\int_{\Gamma_{\ell}}\tilde{\bm{u}}^{j}\cdot\bm{j}$. Find $p^{j}\in Q_{h}$ such that:
$\frac{\tau}{\rho}(\nabla p^{j},\nabla q)_{\Omega_{h}}+\sum_{\ell=1}^{K}\frac{\overline{p^{j}}_{\Gamma_{\ell}}\ \overline{q}_{\Gamma_{\ell}}}{\gamma_{\ell}}+\epsilon\sum_{\ell=1}^{K}(\mathcal{T}(\nabla p^{j}),\mathcal{T}(\nabla q))_{\Gamma_{\ell}}=\sum_{\ell=1}^{K}\bigg{(}\tilde{Q}^{j}+\frac{\alpha_{\ell}\pi_{\ell}^{j-1}}{\gamma_{\ell}}\bigg{)}\overline{q}_{\Gamma_{\ell}}-(\nabla\cdot\tilde{\bm{u}}^{j},q)_{\Omega_{h}},$ (21)
[…]
3. Velocity correction Step: Find $\bm{u}^{j}\in[L^{2}(\Omega_{h})]^{3}$ such that:
$(\bm{u}^{j},\mathbf{v})_{\Omega_{h}}=(\tilde{\bm{u}}^{j}-\frac{\tau}{\rho}\nabla p^{j},\mathbf{v})_{\Omega_{h}}$
@@ Appendix B ROUKF algorithm
Here we detail the ROUKF algorithm adapted from [10]. Let us first consider the notation $[\bm{Z}^{(*)}]$ as the matrix with the column-wise collection of vectors $\bm{Z}_{(1)},\bm{Z}_{(2)},\dots$.
Define the canonical sigma-points $\bm{I}_{(1)},\dots,\bm{I}_{(2p)}\in\mathbb{R}^{p}$ such that
$\bm{I}_{(i)}=\begin{cases}\sqrt{p}\bm{e}_{i},&\text{for }1\leq i\leq p\\ -\sqrt{p}\bm{e}_{i-p},&\text{for }p+1\leq i\leq 2p\end{cases}$
where the vectors $\bm{e}_{i}$ form the canonical base of of $\mathbb{R}^{p}$. Moreover, define the weight $\alpha=\frac{1}{2p}$.
We denote by $\bm{\hat{X}}^{n}_{-},\bm{\hat{X}}^{n}_{+}\in\mathbb{R}^{r}$ a priori (model prediction) and a posteriori (corrected by observations) estimates of the true state $\bm{X}^{n}\in\mathbb{R}^{r}$. In the semi-implicit coupled 3D-0D fractional step Algorithm 1, the state consists in the velocity field $\bm{u}^{n}$ and the Windkessel pressures $\pi^{n}_{\ell}$. Estimates of all unknown parameters are summarized by the corresponding a priori and a posteriori vectors $\bm{\hat{\theta}}^{n}_{-},\bm{\hat{\theta}}^{n}_{+}\in\mathbb{R}^{p}$. The discretized forward model is written as $\bm{X}^{n}=\mathcal{A}^{n}(\bm{X}^{n-1},\bm{\theta}^{n-1})$, $\mathcal{A}^{n}$ denoting the model operator.
For given values of the initial condition $\bm{\hat{X}}^{0}_{+}=\bm{X}^{0}\in\mathbb{R}^{r}$, the initial expected value of the parameters $\bm{\hat{\theta}}^{0}_{+}=\bm{\theta}^{0}\in\mathbb{R}^{p}$ and its covariance matrix $\bm{P}^{0}$, perform
@@
    $\hskip 2.84544pt\bm{L}_{\theta}^{0}=\sqrt{\bm{P}^{0}}\ \text{ (Cholesky factor)},\quad\bm{L}_{X}^{0}=\bm{0}\in\mathbb{R}^{r\times p},\quad\bm{U}_{0}=\bm{P}_{\alpha}\equiv\alpha[\bm{I}^{(*)}][\bm{I}^{(*)}]^{T}$ (22a)
@@
    $\hskip 2.84544pt\begin{cases}\bm{\hat{X}}^{n-1}_{(i)}=\bm{\hat{X}}^{n-1}_{+}+\bm{L}_{X}^{n-1}(\bm{C}^{n-1})^{T}\bm{I}_{(i)},\\ \bm{\hat{\theta}}^{n-1}_{(i)}=\bm{\hat{\theta}}^{n-1}_{+}+\bm{L}^{n-1}_{\theta}(\bm{C}^{n-1})^{T}\bm{I}_{(i)}\end{cases}$ (22b)
    with $\bm{C}^{n-1}$ the Cholesky factor of $(\bm{U}^{n-1})^{-1}$.
@@
    $\hskip 2.84544pt\begin{cases}\bm{\hat{X}}^{n}_{(i)}=\mathcal{A}^{n}(\bm{\hat{X}}^{n-1}_{(i)},\bm{\hat{\theta}}^{n-1}_{(i)}),\quad\bm{\hat{\theta}}^{n}_{(i)}=\bm{\hat{\theta}}^{n-1}_{(i)},\quad i=1,\dots,2p\\ \bm{\hat{X}}^{n}_{-}=E_{\alpha}([(\bm{\hat{X}}^{n})^{(*)}])\equiv\alpha\sum_{i=1}^{2p}\bm{\hat{X}}^{n}_{(i)}\\ \bm{\hat{\theta}}^{n}_{-}=E_{\alpha}([(\bm{\hat{\theta}}^{n})^{(*)}])\end{cases}$ (22c)
  • Correction: compute a posteriori estimates based on measurements for state and parameters, using the $i$-th particle innovation $\bm{\Gamma}^{n}_{(i)}$:
    $\hskip 2.84544pt\begin{cases}\bm{L}^{n}_{X}=\alpha[(\bm{\hat{X}}^{n})^{(*)}][\bm{I}^{(*)}]^{T}\\ \bm{L}_{\theta}^{n}=\alpha[(\bm{\hat{\theta}}^{n})^{(*)}][\bm{I}^{(*)}]^{T}\\ \bm{L}_{\Gamma}^{n}=\alpha[(\bm{\Gamma}^{n})^{(*)}][\bm{I}^{(*)}]^{T}\\ \bm{U}^{n}=\bm{P}_{\alpha}+(\bm{L}^{n}_{\Gamma})^{T}\bm{W}^{-1}\bm{L}^{n}_{\Gamma}\ ,\quad\bm{P}_{\alpha}=\alpha[\bm{I}^{(*)}][\bm{I}^{(*)}]^{T}\\ \bm{\hat{X}}^{n}_{+}=\bm{\hat{X}}^{n}_{-}-\bm{L}^{n}_{X}(\bm{U}^{n})^{-1}(\bm{L}^{n}_{\Gamma})^{T}\bm{W}^{-1}E_{\alpha}([(\bm{\Gamma}^{n})^{(*)}])\\ \bm{\hat{\theta}}_{+}^{n}=\bm{\hat{\theta}}^{n}_{-}-\bm{L}^{n}_{\theta}(\bm{U}^{n})^{-1}(\bm{L}^{n}_{\Gamma})^{T}\bm{W}^{-1}E_{\alpha}([(\bm{\Gamma}^{n})^{(*)}])\\ \bm{P}_{\theta}^{n}=\bm{L}_{\theta}^{n}(\bm{U}^{n})^{-1}(\bm{L}_{\theta}^{n})^{T}\\ \bm{P}_{X}^{n}=\bm{L}_{X}^{n}(\bm{U}^{n})^{-1}(\bm{L}_{X}^{n})^{T}\end{cases}$ (22d)
with $\bm{W}=\sigma_{y}^{2}\mathbb{I}$.

### v2

@@ 6 Conclusion
We have proposed a new formulation for the inverse problem for parameter estimation in fluid flow problems using undersampled frequency-space MRI data and demonstrated it using a Reduced-Order Unscented Kalman Filter. This method outperforms the results of parameter estimation using velocity data reconstructed with Compressed Sensing, especially for high undersampling rates and different vencvenc values.
@@ Acknowledgments
C.B. and M.L. acknowledge the funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No 852544 - CardioZoom).
@@ Appendix A Numerical solution method of the forward problem
Given the initial conditions $\boldsymbol{u}^{0}=\boldsymbol{u}(0)\in V_{\Gamma_{w},h}$ and $\pi^{0}_{1},\dots,\pi_{N}^{0}\in\mathbb{R}$, perform for $j>0$, with $t^{j}=j\tau$:
1. Viscous Step: Find the tentative velocity $\tilde{\boldsymbol{u}}^{n}\in V_{\Gamma_{w},h}$ such that:
$\begin{cases}\displaystyle\tilde{\boldsymbol{u}}^{j}|_{\Gamma_{in}}=\boldsymbol{u}_{inlet}(t^{j})\\ \displaystyle\frac{\tau}{\rho}(\tilde{\boldsymbol{u}}^{j},\boldsymbol{v})_{\Omega_{h}}+\rho(\boldsymbol{u}^{j-1}\cdot\nabla\tilde{\boldsymbol{u}}^{j},\boldsymbol{v})_{\Omega_{h}}+\frac{\rho}{2}((\nabla\cdot\boldsymbol{u}^{j-1})\tilde{\boldsymbol{u}}^{j},\boldsymbol{v})_{\Omega_{h}}+(\delta\boldsymbol{u}^{j-1}\cdot\nabla\tilde{\boldsymbol{u}}^{j},\boldsymbol{u}^{j-1}\cdot\nabla\boldsymbol{v})_{\Omega_{h}}\\ \displaystyle+2\mu(\epsilon(\tilde{\boldsymbol{u}}^{j}),\epsilon(\boldsymbol{v}))_{\Omega_{h}}+\sum_{\ell=1}^{K}\frac{\rho}{2}|\boldsymbol{u}^{j-1}\cdot\mathbf{n}|_{-}(\tilde{\boldsymbol{u}}^{j},\boldsymbol{v})_{\Gamma_{\ell}}=\frac{\tau}{\rho}(\boldsymbol{u}^{j-1},\boldsymbol{v})_{\Omega_{h}}\end{cases}$ (21)
@@ 1. Viscous Step: Find the tentative velocity $\tilde{\boldsymbol{u}}^{n}\in V_{\Gamma_{w},h}$ such that:
2. Projection-Windkessel Step: Compute $\tilde{Q}^{j}=\int_{\Gamma_{\ell}}\tilde{\boldsymbol{u}}^{j}\cdot\boldsymbol{j}$. Find $p^{j}\in Q_{h}$ such that:
$\frac{\tau}{\rho}(\nabla p^{j},\nabla q)_{\Omega_{h}}+\sum_{\ell=1}^{K}\frac{\overline{p^{j}}_{\Gamma_{\ell}}\ \overline{q}_{\Gamma_{\ell}}}{\gamma_{\ell}}+\epsilon\sum_{\ell=1}^{K}(\mathcal{T}(\nabla p^{j}),\mathcal{T}(\nabla q))_{\Gamma_{\ell}}=\sum_{\ell=1}^{K}\bigg(\tilde{Q}^{j}+\frac{\alpha_{\ell}\pi_{\ell}^{j-1}}{\gamma_{\ell}}\bigg)\overline{q}_{\Gamma_{\ell}}-(\nabla\cdot\tilde{\boldsymbol{u}}^{j},q)_{\Omega_{h}},$ (22)
for all $q\in Q_{h}$ and with $\overline{(\cdot)}_{\Gamma_{\ell}}=\frac{1}{Area(\Gamma_{\ell})}\int_{\Gamma_{\ell}}(\cdot)ds$ and $\mathcal{T}(\mathbf{f})=\mathbf{f}-(\mathbf{f}\cdot\mathbf{j})\mathbf{j}$.
3. Velocity correction Step: Find $\boldsymbol{u}^{j}\in[L^{2}(\Omega_{h})]^{3}$ such that:
$(\boldsymbol{u}^{j},\mathbf{v})_{\Omega_{h}}=(\tilde{\boldsymbol{u}}^{j}-\frac{\tau}{\rho}\nabla p^{j},\mathbf{v})_{\Omega_{h}}$
@@ Appendix B ROUKF algorithm
Here we detail the ROUKF algorithm adapted from [moireau-chapelle-11]. Let us first consider the notation $\boldsymbol{Z}^{(*)}]$ as the matrix with the column-wise collection of vectors $\boldsymbol{Z}_{(1)},\boldsymbol{Z}_{(2)},\dots$.
Define the canonical sigma-points $\boldsymbol{I}_{(1)},\dots,\boldsymbol{I}_{(2p)}\in\mathbb{R}^{p}$ such that
$\boldsymbol{I}_{(i)}=\begin{cases}\sqrt{p}\boldsymbol{e}_{i},&\text{for }1\leq i\leq p\\ -\sqrt{p}\boldsymbol{e}_{i-p},&\text{for }p+1\leq i\leq 2p\end{cases}$
where the vectors $\boldsymbol{e}_{i}$ form the canonical base of of $\mathbb{R}^{p}$. Moreover, define the weight $\alpha=\frac{1}{2p}$.
We denote by $\boldsymbol{\hat{X}}^{n}_{-},\boldsymbol{\hat{X}}^{n}_{+}\in\mathbb{R}^{r}$ a priori (model prediction) and a posteriori (corrected by observations) estimates of the true state $\boldsymbol{X}^{n}\in\mathbb{R}^{r}$. In the semi-implicit coupled 3D-0D fractional step Algorithm 1, the state consists in the velocity field $\boldsymbol{u}^{n}$ and the Windkessel pressures $\pi^{n}_{\ell}$. Estimates of all unknown parameters are summarized by the corresponding a priori and a posteriori vectors $\boldsymbol{\hat{\theta}}^{n}_{-},\boldsymbol{\hat{\theta}}^{n}_{+}\in\mathbb{R}^{p}$. The discretized forward model is written as $\boldsymbol{X}^{n}=\mathcal{A}^{n}(\boldsymbol{X}^{n-1},\boldsymbol{\theta}^{n-1})$, $\mathcal{A}^{n}$ denoting the model operator.
For given values of the initial condition $\boldsymbol{\hat{X}}^{0}_{+}=\boldsymbol{X}^{0}\in\mathbb{R}^{r}$, the initial expected value of the parameters $\boldsymbol{\hat{\theta}}^{0}_{+}=\boldsymbol{\theta}^{0}\in\mathbb{R}^{p}$ and its covariance matrix $\boldsymbol{P}^{0}$, perform
@@
    $\boldsymbol{L}_{\theta}^{0}=\sqrt{\boldsymbol{P}^{0}}\ \text{ (Cholesky factor)},\boldsymbol{L}_{X}^{0}=\boldsymbol{0}\in\mathbb{R}^{r\times p},\boldsymbol{U}_{0}=\boldsymbol{P}_{\alpha}\equiv\alpha[\boldsymbol{I}^{(*)}][\boldsymbol{I}^{(*)}]^{T}$ (23a)
@@
    $\hskip 2.84544pt\begin{cases}\boldsymbol{\hat{X}}^{n-1}_{(i)}=\boldsymbol{\hat{X}}^{n-1}_{+}+\boldsymbol{L}_{X}^{n-1}(\boldsymbol{C}^{n-1})^{T}\boldsymbol{I}_{(i)},\\ \boldsymbol{\hat{\theta}}^{n-1}_{(i)}=\boldsymbol{\hat{\theta}}^{n-1}_{+}+\boldsymbol{L}^{n-1}_{\theta}(\boldsymbol{C}^{n-1})^{T}\boldsymbol{I}_{(i)}\end{cases}$ (23b)
    with $\boldsymbol{C}^{n-1}$ the Cholesky factor of $(\boldsymbol{U}^{n-1})^{-1}$.
@@
    $\hskip 2.84544pt\begin{cases}\boldsymbol{\hat{X}}^{n}_{(i)}=\mathcal{A}^{n}(\boldsymbol{\hat{X}}^{n-1}_{(i)},\boldsymbol{\hat{\theta}}^{n-1}_{(i)}),\quad\boldsymbol{\hat{\theta}}^{n}_{(i)}=\boldsymbol{\hat{\theta}}^{n-1}_{(i)},\quad i=1,\dots,2p\\ \boldsymbol{\hat{X}}^{n}_{-}=E_{\alpha}([(\boldsymbol{\hat{X}}^{n})^{(*)}])\equiv\alpha\sum_{i=1}^{2p}\boldsymbol{\hat{X}}^{n}_{(i)}\\ \boldsymbol{\hat{\theta}}^{n}_{-}=E_{\alpha}([(\boldsymbol{\hat{\theta}}^{n})^{(*)}])\end{cases}$ (23c)
  • Correction: compute a posteriori estimates based on measurements for state and parameters, using the $i$-th particle innovation $\boldsymbol{\Gamma}^{n}_{(i)}$:
    $\hskip 2.84544pt\begin{cases}\boldsymbol{L}^{n}_{X}=\alpha[(\boldsymbol{\hat{X}}^{n})^{(*)}][\boldsymbol{I}^{(*)}]^{T}\\ \boldsymbol{L}_{\theta}^{n}=\alpha[(\boldsymbol{\hat{\theta}}^{n})^{(*)}][\boldsymbol{I}^{(*)}]^{T}\\ \boldsymbol{L}_{\Gamma}^{n}=\alpha[(\boldsymbol{\Gamma}^{n})^{(*)}][\boldsymbol{I}^{(*)}]^{T}\\ \boldsymbol{U}^{n}=\boldsymbol{P}_{\alpha}+(\boldsymbol{L}^{n}_{\Gamma})^{T}\boldsymbol{W}^{-1}\boldsymbol{L}^{n}_{\Gamma}\ ,\quad\boldsymbol{P}_{\alpha}=\alpha[\boldsymbol{I}^{(*)}][\boldsymbol{I}^{(*)}]^{T}\\ \boldsymbol{\hat{X}}^{n}_{+}=\boldsymbol{\hat{X}}^{n}_{-}-\boldsymbol{L}^{n}_{X}(\boldsymbol{U}^{n})^{-1}(\boldsymbol{L}^{n}_{\Gamma})^{T}\boldsymbol{W}^{-1}E_{\alpha}([(\boldsymbol{\Gamma}^{n})^{(*)}])\\ \boldsymbol{\hat{\theta}}_{+}^{n}=\boldsymbol{\hat{\theta}}^{n}_{-}-\boldsymbol{L}^{n}_{\theta}(\boldsymbol{U}^{n})^{-1}(\boldsymbol{L}^{n}_{\Gamma})^{T}\boldsymbol{W}^{-1}E_{\alpha}([(\boldsymbol{\Gamma}^{n})^{(*)}])\\ \boldsymbol{P}_{\theta}^{n}=\boldsymbol{L}_{\theta}^{n}(\boldsymbol{U}^{n})^{-1}(\boldsymbol{L}_{\theta}^{n})^{T}\\ \boldsymbol{P}_{X}^{n}=\boldsymbol{L}_{X}^{n}(\boldsymbol{U}^{n})^{-1}(\boldsymbol{L}_{X}^{n})^{T}\end{cases}$ (23d)
with $\boldsymbol{W}=\sigma_{y}^{2}\mathbb{I}$.
Ethics Statement
None